A friend lends you $200 for a week, which you agree to repay with 7% one time interest. How much will you have to repay?
Answer #1

Amount lended=$200

Interest rate=7%

Amount to repay=200+7% of 200

=200+(7×200)/100

=200+14

=$214

Amount to repay=$214